# Elastic-plastic cell-based smoothed finite element method solving geotechnical problems

### v2

@@ Abstract
An elastic–plastic cell-based smoothed finite element method (CSFEM) is proposed for geotechnical analysis of soils and rocks exhibiting nonlinear and path-dependent behaviors. By introducing strain smoothing over subcell domains and employing a consistent stress return-mapping algorithm, the method enhances stress accuracy, alleviates volumetric locking, and reduces sensitivity to mesh distortion while retaining the flexibility of polygonal elements. The formulation is implemented in ABAQUS via a user-defined element and validated through benchmark and practical problems, including a pressurized thick cylinder, biaxial soil test, strip footing bearing capacity, tunnel excavation, and slope stability. Numerical results show excellent agreement with analytical solutions and conventional FEM, with smoother stress fields, improved convergence, and higher accuracy in ultimate load prediction. These findings demonstrate that CSFEM provides a stable and efficient framework for elastic–plastic analysis of complex geotechnical problems.
@@
cell-based smoothed finite element method (CSFEM), elastic–plastic analysis, geotechnical engineering, ABAQUS UEL implementation
@@
[inst1]organization=Department One,addressline=Address One, city=City One, postcode=00000, state=State One, country=Country One
@@
[inst2]organization=Department Two,addressline=Address Two, city=City Two, postcode=22222, state=State Two, country=Country Two
@@ 1 Introduction
The reliable prediction of the mechanical response of geomaterials under complex loading conditions remains a fundamental challenge in geotechnical engineering. Geomaterials, such as soil and rock, often exhibit nonlinear behaviors including plasticity, strain-softening, and path-dependent responses (Chu et al., 1992; Kok et al., 2009). These features require robust numerical methods to capture the intricate interplay between elasticity and plasticity in large-scale simulations of foundations (Ray et al., 2021), slopes (Kardani et al., 2021), tunnels (Jin et al., 2022), embankments(Xia et al., 2024), and underground structures(Hemeda, 2022). Among various computational approaches, the finite element method (FEM) has become the most widely adopted due to its flexibility in handling irregular geometries and boundary conditions (Bottero et al., 1980; Liu et al., 2015; Wei et al., 2025).
@@
However, FEM formulations present several drawbacks in the elastic–plastic analysis of geotechnical problems. FEM is known to overestimate the stiffness matrix, leading to an overall response that is too stiff and an underestimation of the internal strain energy (Fraeijs de Veubeke, 1965). As a result, the computed displacements are generally smaller than the actual values. In addition, the required mapping and coordinate transformation restrict element shapes: for a four-node isoparametric element, all interior angles must remain below 180 and the Jacobian determinant must remain positive (Liu and Quek, ; Liu et al., ; Liu et al., 2007). These constraints increase computational cost and significantly limit FEM performance in geotechnical problems involving large deformation or severe mesh distortion.
@@
By introducing the strain-smoothing technique first proposed by Liu et al. (Liu et al., 2007), the smoothed finite element method (SFEM) (Liu and Trung, 2016) has developed into a promising framework for overcoming the intrinsic limitations of classical FEM, thereby drawing substantial attention in recent years. The SFEM delivers higher accuracy and faster energy convergence than the conventional FEM (Dai et al., 2007; Yan et al., 2025). Specifically, its freedom from mapping or coordinate transformation allows elements of arbitrary shapes, making it particularly beneficial for geotechnical analyses where large deformations can cause severe mesh distortion.
@@
Several SFEM variants have been proposed, including face-based (Huang et al., 2023; Jiang et al., 2025), node-based (Sun et al., 2022; Lyu et al., 2024), edge-based (He, 2022; He et al., 2022), and cell-based formulations (Surendran et al., 2021; Liu et al., 2022; Sun et al., 2025).Among these, the cell-based SFEM (CSFEM) provides a particularly straightforward framework that remains close to the standard FEM procedure. In the CSFEM, each element is subdivided into smoothing sub-cells where the strain field is averaged, and the stiffness contribution is evaluated through surface integration based on the divergence theorem (Lee et al., 2015).
@@
At present, the CSFEM has been successfully applied to several physical field problems: Surendran et al.(Surendran et al., 2021) presents a CSFEM using polygonal elements to model interfacial cracks with non-matching grids, enabling accurate fracture analysis while reducing degrees of freedom and relaxing meshing constraints. Yan et al. (Yan et al., 2025) present a fast CSFEM to solve static and dynamic problems. Lee et al. (Lee et al., 2015) a three-dimensional CSFEM for application to elastic-plastic analysis for the solid mechanics. Peng et al. (Peng et al., 2021) developed a novel phase-field model for brittle fracture using CSFEM with spectral decomposition to capture nonlinear stress and elastic responses. These studies demonstrate the effectiveness and versatility of CSFEM, yet its application in geotechnical engineering simulations remains limited.
@@
In this work, we propose an elastic–plastic CSFEM tailored for geotechnical applications. The method incorporates smoothing operations into standard FEM formulations. The proposed framework enhances numerical stability, reduces sensitivity to mesh distortion, and improves the accuracy of stress prediction in nonlinear problems. The remainder of this paper is organized as follows. Section 2 outlines the theoretical framework of the proposed elastic-plastic CSFEM, including the smoothing formulation and elastic–plastic integration scheme. Section 5 details the numerical implementation and integration of elastic-plastic CSFEM into ABAQUS via a user-defined element (UEL). Section 6 presents benchmark numerical studies to verify accuracy and convergence. Finally, Section 7 concludes with key findings and future research directions.
2 Cell-based SFEM formulation for elasto-plasticity
@@
In this study, we examine a deformable body that occupies the domain $\Omega$. This body responds to internal body forces b, externally applied tractions $\overline{\textbf{t}}$ on the boundary $\Gamma_{t}$, and displacement boundary conditions $\mathbf{u}=\overline{\mathbf{u}}$ on $\Gamma_{u}$, as depicted in Figure 1. The governing equation and boundary conditions are expressed as follows:
@@
$\boldsymbol{\sigma}=\boldsymbol{D}\boldsymbol{\varepsilon},$ (2)
where $\boldsymbol{D}$ is the constitutive matrix.
@@
where $\mathbf{d}_{i}=\left[\begin{array}[]{ll}u_{i}&v_{i}\end{array}\right]^{\mathrm{T}}$ is the nodal displacement vector, and $\boldsymbol{B}_{i}$ is the strain-displacement matrix as follows:
@@
$\int_{\Omega}\delta\mathbf{d}^{\mathrm{T}}\mathbf{B}^{\mathrm{T}}\boldsymbol{\sigma}\mathrm{d}\Omega-\int_{\Omega}\delta\mathbf{d}^{\mathrm{T}}\mathbf{N}^{\mathrm{T}}[\mathbf{b}-\rho\ddot{\mathbf{u}}-c\mathbf{u}]\mathrm{d}\Omega-\int_{\Gamma_{t}}\delta\mathbf{d}^{\mathrm{T}}\mathbf{N}^{\mathrm{T}}\mathbf{t}\mathrm{d}\Gamma=0,$ (6)
@@
$\mathbf{Kd}=\mathbf{P}.$ (8)
$\mathbf{K}=\int_{\Omega}\mathbf{B}^{T}\mathbf{DB}\mathrm{d}\Omega,$ (9)
$\mathbf{P}=\int_{\Omega}\mathbf{N}^{T}\mathbf{b}\mathrm{d}\Omega+\int_{\Gamma_{t}}\mathbf{N}^{T}\mathbf{t}\mathrm{d}\Gamma,$ (10)
where $\mathbf{K}$ is the stiffness, mass, and damping matrices respectively; $\mathbf{P}$ is the external force vector.
The modified strain field $\tilde{\boldsymbol{\varepsilon}}^{h}$ is determined through a weighted-average formulation of the conventional strain field
@@
where $A_{C}$ is the area of the sub-cell. The modified strain field $\tilde{\boldsymbol{\varepsilon}}^{h}$ is computed from the smoothed discretized gradient operator $\widetilde{\mathbf{B}}$ and the nodal displacement $\mathbf{b}$. By applying the technique of integration by parts to Eq.(11), we obtain:
$\tilde{\varepsilon}^{h}\left(\mathbf{x}_{C}\right)=\int_{\Gamma}\mathbf{u}^{h}(\mathbf{x})\mathrm{n}(\mathbf{x})\Phi\left(\mathbf{x}-\mathbf{x}_{C}\right)d\Gamma-\int_{\Omega}\mathbf{u}^{h}(\mathbf{x})\nabla\Phi\left(\mathbf{x}-\mathbf{x}_{C}\right)d\Omega.$ (13)
@@
$\tilde{\boldsymbol{\varepsilon}}^{h}\left(\mathbf{x}_{C}\right)=\int_{\Gamma_{C}}\mathbf{u}^{h}(x)\mathbf{n}(x)\Phi\left(\mathbf{x}-\mathbf{x}_{C}\right)d\Gamma=\frac{1}{A_{c}}\int_{\Gamma_{c}}\mathbf{u}^{h}(\mathbf{x})\mathbf{n}(\mathbf{x})d\Gamma,$ (14)
@@ 3 Elasto-plastic constitutive formulation
In elasto-plastic analysis, the total incremental strain is conventionally decomposed into its elastic and plastic components as
$\boldsymbol{\Delta\varepsilon}=\boldsymbol{\Delta\varepsilon}^{\mathrm{e}}+\boldsymbol{\Delta\varepsilon}^{\mathrm{p}},$ (19)
where $\boldsymbol{\Delta\varepsilon}^{\mathrm{e}}$ and $\boldsymbol{\Delta\varepsilon}^{\mathrm{p}}$ denote the elastic and plastic strain increments, respectively. The incremental stress–strain relationship for the elastic response follows Hooke’s law, given by
$\boldsymbol{\Delta\sigma}=\mathbf{D}:\left(\boldsymbol{\Delta\varepsilon}-\boldsymbol{\Delta\varepsilon}^{\mathrm{p}}\right)=\mathbf{D}:\boldsymbol{\Delta\varepsilon}^{\mathrm{e}},$ (20)
@@
$\boldsymbol{\Delta\varepsilon}^{\mathrm{p}}=\frac{\partial g}{\partial\boldsymbol{\sigma}}\,\Delta\lambda=\frac{\partial F}{\partial\boldsymbol{\sigma}}\,\Delta\lambda,$ (21)
@@
$\Delta F=\frac{\partial F}{\partial\boldsymbol{\sigma}}:\boldsymbol{\Delta\sigma}+\frac{\partial F}{\partial\kappa}\,\Delta\kappa=0,$ (22)
where $\kappa$ denotes the hardening parameter. The evolution of $\kappa$ is described by the hardening law
$\Delta\kappa=H\,\Delta\lambda,$ (23)
with $H$ being the plastic hardening modulus.
In this study, the Mohr–Coulomb failure criterion is adopted to assess the stability of vegetated slopes. The yield function is defined as
$F=\left(\sqrt{3}\cos\theta_{0}-\sin\theta_{0}\sin\phi\right)q-3p\sin\phi-3(c_{r}+c_{s})\cos\phi=0,$ (24)
where $p=\tfrac{1}{3}(\sigma_{1}+\sigma_{2}+\sigma_{3})$ is the mean stress, $q=\sqrt{\tfrac{3}{2}\boldsymbol{s}:\boldsymbol{s}}$ is the equivalent deviatoric stress, $\theta_{0}$ is the Lode angle, $\phi$ is the internal friction angle, and $c_{r}$ and $c_{s}$ denote the root and soil cohesion, respectively.
By combining the flow rule, the consistency condition, and the hardening law, the elasto-plastic constitutive matrix can be derived as
$\mathbf{D}_{ep}=\mathbf{D}-\mathbf{D}\left(\frac{\partial F}{\partial\boldsymbol{\sigma}}\right)\left(\frac{\partial F}{\partial\boldsymbol{\sigma}}\right)^{T}\mathbf{D}\left(A+\left(\frac{\partial F}{\partial\boldsymbol{\sigma}}\right)^{T}\mathbf{D}\left(\frac{\partial F}{\partial\boldsymbol{\sigma}}\right)\right)^{-1},$ (25)
where $A=-\dfrac{1}{\Delta\lambda}\dfrac{\partial F}{\partial\kappa}\Delta\kappa$ and $\Delta\lambda$ is the plastic multiplier. The elasto-plastic constitutive matrix $\mathbf{D}_{ep}$ provides the incremental stress–strain relationship during yielding and varies spatially within the domain, reflecting the nonlinear mechanical response of the material.

### v3

# Elasto-plastic cell-based smoothed finite element method solving geotechnical problems

@@ Abstract
This work develops an elasto-plastic cell-based smoothed finite element method (CSFEM) for geotechnical analysis. The formulation incorporates a smoothed strain field into the standard elasto-plastic framework based on the Mohr–Coulomb criterion and is implemented in ABAQUS through a user-defined element (UEL). A UEL–UMAT data-transfer strategy is introduced to enable post-processing of stress and strain in ABAQUS. The method is assessed using several benchmark problems, including three classical examples, a tunnel excavation, and a slope stability analysis. The results show that the CSFEM achieves accuracy comparable to or slightly better than the conventional FEM and matches analytical or reference solutions for the examined cases. These findings indicate that the proposed CSFEM provides a practical and robust alternative for routine elasto-plastic analyses in geotechnical engineering.
@@
Cell-based smoothed finite element method (CSFEM), Elasto-plastic analysis, Geotechnical engineering, ABAQUS UEL implementation
@@
{highlights}
Elasto-plastic CSFEM for geotechnical analysis is developed.
Implement the method in ABAQUS via a UEL with a UEL–UMAT transfer scheme.
Benchmarks show accuracy comparable to FEM with smoother contours.
@@ 1 Introduction
The reliable prediction of the mechanical response of geomaterials under complex loading conditions remains a fundamental challenge in geotechnical engineering. Geomaterials, such as soil and rock, often exhibit nonlinear behaviors including plasticity, strain-softening, and path-dependent responses (chu1992strain; kok2009case). These features require robust numerical methods to capture the intricate interplay between elasticity and plasticity in large-scale simulations of foundations (ray2021application), slopes (kardani2021improved), tunnels (jin2022effect), embankments(xia2024stability), and underground structures(hemeda2022geotechnical). Among various computational approaches, the finite element method (FEM) has become the most widely adopted due to its flexibility in handling irregular geometries and boundary conditions (bottero1980finite; liu2015slope; wei20253d).
@@
However, FEM formulations present several drawbacks in the elasto-plastic analysis of geotechnical problems. FEM is known to overestimate the stiffness matrix, leading to an overall response that is too stiff and an underestimation of the internal strain energy (fraeijs1965displacement). As a result, the computed displacements are generally smaller than the actual values. In addition, the required mapping and coordinate transformation restrict element shapes: for a four-node isoparametric element, all interior angles must remain below 180 and the Jacobian determinant must remain positive (liuFiniteElementMethod2003; liuTheoreticalAspectsSmoothed2007; liuSmoothedFiniteElement2007). These constraints increase computational cost and significantly limit FEM performance in geotechnical problems involving large deformation or severe mesh distortion.
@@
By introducing the strain-smoothing technique first proposed by Liu et al. (liuSmoothedFiniteElement2007), the smoothed finite element method (SFEM) (liu2016smoothed) has developed into a promising framework for overcoming the intrinsic limitations of classical FEM, thereby drawing substantial attention in recent years. The SFEM delivers higher accuracy and faster energy convergence than the conventional FEM (daiNsidedPolygonalSmoothed2007; yanFastCellSmoothedFinite2025). Specifically, its freedom from mapping or coordinate transformation allows elements of arbitrary shapes, making it particularly beneficial for geotechnical analyses where large deformations can cause severe mesh distortion.
@@
Several SFEM variants have been proposed, including face-based (huang2023face; jiang2025strength), node-based (sun2022stable; lyu2024implicit), edge-based (he2022edgeon; he2022edgean), and cell-based formulations (surendran2021cell; liu2022cell; sun2025cell).Among these, the cell-based SFEM (CSFEM) provides a particularly straightforward framework that remains close to the standard FEM procedure. In the CSFEM, each element is subdivided into smoothing sub-cells where the strain field is averaged, and the stiffness contribution is evaluated through surface integration based on the divergence theorem (lee2015three).
@@
At present, the CSFEM has been successfully applied to several physical field problems: Surendran et al.(surendran2021cell) presents a CSFEM using polygonal elements to model interfacial cracks with non-matching grids, enabling accurate fracture analysis while reducing degrees of freedom and relaxing meshing constraints. Yan et al. (yanFastCellSmoothedFinite2025) present a fast CSFEM to solve static and dynamic problems. Lee et al. (lee2015three) a three-dimensional CSFEM for application to elasto-plastic analysis for the solid mechanics. Peng et al. (peng2021phase) developed a novel phase-field model for brittle fracture using CSFEM with spectral decomposition to capture nonlinear stress and elastic responses. These studies demonstrate the effectiveness and versatility of CSFEM, yet its application in geotechnical engineering simulations remains limited.
@@
In this work, we propose an elasto-plastic CSFEM tailored for geotechnical applications. The method incorporates strain smoothing into the standard FEM formulation, replacing the compatible strain with a smoothed strain field evaluated over subcells and coupling it with a Mohr–Coulomb elasto-plastic model. This leads to a softer and more stable stiffness matrix, reduces mesh-distortion sensitivity, and improves stress prediction in nonlinear analyses. The remainder of this paper is organized as follows. Section 2 presents the theoretical formulation of the elasto-plastic CSFEM. Section 3 introduces the Mohr–Coulomb constitutive model. Section 4 describes the construction of smoothing subcells and the smoothed strain–displacement matrix. Section 5 outlines the ABAQUS UEL implementation and the UEL–UMAT data-transfer scheme. Section 6 provides benchmark examples to assess accuracy and convergence. Finally, Section 7 summarizes the main findings.
2 CSFEM formulation for elasto-plasticity
@@
In this study, we examine a deformable body that occupies the domain $\Omega$. This body responds to internal body forces b, externally applied tractions $\overline{\textbf{t}}$ on the boundary $\Gamma_{t}$, and displacement boundary conditions $\mathbf{u}=\overline{\mathbf{u}}$ on $\Gamma_{u}$. The governing equation and boundary conditions are expressed as follows:
@@
$\boldsymbol{\sigma}=\boldsymbol{\mathrm{D}}\boldsymbol{\varepsilon},$ (2)
where $\boldsymbol{\mathrm{D}}$ is the constitutive matrix.
@@
where $\mathbf{d}_{i}=\left[\begin{array}[]{ll}u_{i}&v_{i}\end{array}\right]^{\mathrm{T}}$ is the nodal displacement vector, and $\mathbf{B}_{i}$ is the strain-displacement matrix as follows:
@@
$\int_{\Omega}\delta\mathbf{d}^{\mathrm{T}}\mathbf{B}^{\mathrm{T}}\boldsymbol{\sigma}\mathrm{d}\Omega-\int_{\Omega}\delta\mathbf{d}^{\mathrm{T}}\mathbf{N}^{\mathrm{T}}\mathbf{b}\mathrm{d}\Omega-\int_{\Gamma_{t}}\delta\mathbf{d}^{\mathrm{T}}\mathbf{N}^{\mathrm{T}}\mathbf{t}\mathrm{d}\Gamma=0.$ (6)
@@
$\mathbf{Kd}=\mathbf{P},$ (8)
$\mathbf{K}=\int_{\Omega}\mathbf{B}^{\mathrm{T}}\mathbf{DB}\mathrm{d}\Omega,$ (9)
$\mathbf{P}=\int_{\Omega}\mathbf{N}^{\mathrm{T}}\mathbf{b}\mathrm{d}\Omega+\int_{\Gamma_{t}}\mathbf{N}^{\mathrm{T}}\mathbf{t}\mathrm{d}\Gamma,$ (10)
where $\mathbf{K}$ is the stiffness matrices respectively; $\mathbf{P}$ is the external force vector.
The modified strain field $\tilde{\boldsymbol{\varepsilon}}^{h}$ is determined through a weighted-average formulation of the conventional strain field (liuSmoothedFiniteElement2007)
@@
where $A_{C}$ is the area of the sub-cell. The modified strain field $\tilde{\boldsymbol{\varepsilon}}^{h}$ is computed from the smoothed discretized gradient operator $\widetilde{\mathbf{B}}$ and the nodal displacement $\mathbf{d}$. By applying the technique of integration by parts to Eq.(11), we obtain:
$\tilde{\varepsilon}^{h}\left(\mathbf{x}_{C}\right)=\int_{\Gamma}\mathbf{u}^{h}(\mathbf{x})\mathrm{n}(\mathbf{x})\Phi\left(\mathbf{x}-\mathbf{x}_{C}\right)\mathrm{d}\Gamma-\int_{\Omega}\mathbf{u}^{h}(\mathbf{x})\nabla\Phi\left(\mathbf{x}-\mathbf{x}_{C}\right)\mathrm{d}\Omega.$ (13)
@@
$\tilde{\boldsymbol{\varepsilon}}^{h}\left(\mathbf{x}_{C}\right)=\int_{\Gamma_{C}}\mathbf{u}^{h}(x)\mathbf{n}(x)\Phi\left(\mathbf{x}-\mathbf{x}_{C}\right)\mathrm{d}\Gamma=\frac{1}{A_{c}}\int_{\Gamma_{c}}\mathbf{u}^{h}(\mathbf{x})\mathbf{n}(\mathbf{x})d\Gamma,$ (14)
@@ 3 Elasto-plastic constitutive formulation
In elasto-plastic analysis, the total incremental strain $\Delta\boldsymbol{\varepsilon}$ is conventionally decomposed into its elastic and plastic components as
$\Delta\boldsymbol{\varepsilon}=\Delta\boldsymbol{\varepsilon}^{\mathrm{e}}+\Delta\boldsymbol{\varepsilon}^{\mathrm{p}},$ (19)
where $\Delta\boldsymbol{\varepsilon}^{\mathrm{e}}$ and $\Delta\boldsymbol{\varepsilon}^{\mathrm{p}}$ denote the elastic and plastic strain increments, respectively. The incremental stress–strain relationship for the elastic response follows Hooke’s law, given by
$\Delta\boldsymbol{\sigma}=\mathbf{D}\left({\Delta\boldsymbol{\varepsilon}}-\Delta\boldsymbol{\varepsilon}^{\mathrm{p}}\right)=\mathbf{D}\Delta\boldsymbol{\varepsilon}^{\mathrm{e}},$ (20)
@@
$\Delta\boldsymbol{\varepsilon}^{\mathrm{p}}=\frac{\partial g}{\partial\boldsymbol{\sigma}}\,\Delta\lambda=\frac{\partial F}{\partial\boldsymbol{\sigma}}\,\Delta\lambda,$ (21)
@@
$\Delta F=\frac{\partial F}{\partial\boldsymbol{\sigma}}\Delta\boldsymbol{\sigma}=0.$ (22)
In this study, the Mohr–Coulomb failure criterion is employed. The yield function is defined as (sherif2012matlab)
$F=\left(\sqrt{3}\cos\theta-\sin\theta\sin\phi\right)q-3p\sin\phi-3c\cos\phi=0,$ (23)
where $p=\tfrac{1}{3}(\sigma_{1}+\sigma_{2}+\sigma_{3})$ is the mean stress, $q=\sqrt{\tfrac{3}{2}\boldsymbol{s}:\boldsymbol{s}}$ is the equivalent deviatoric stress, $\theta$ is the Lode angle, $\phi$ is the internal friction angle, and $c$ denote the soil cohesion, respectively.
By combining the flow rule, the consistency condition, and the hardening law, the elasto-plastic constitutive matrix can be derived as (jabbarilakNonlinearScaledBoundary2019)
$\mathbf{D}_{ep}=\mathbf{D}-\mathbf{D}\left(\frac{\partial F}{\partial\boldsymbol{\sigma}}\right)\left(\frac{\partial F}{\partial\boldsymbol{\sigma}}\right)^{\mathrm{T}}\mathbf{D}\left(\left(\frac{\partial F}{\partial\boldsymbol{\sigma}}\right)^{\mathrm{T}}\mathbf{D}\left(\frac{\partial F}{\partial\boldsymbol{\sigma}}\right)\right)^{-1}.$ (24)
The elasto-plastic constitutive matrix $\mathbf{D}_{ep}$ provides the incremental stress–strain relationship during yielding and varies spatially within the domain, reflecting the nonlinear mechanical response of the material.
In the numerical solution procedure, the equilibrium equation is expressed as
$\mathbf{R}=\mathbf{F}_{\mathrm{ext}}-\mathbf{F}_{\mathrm{int}}=\mathbf{0},$ (25)
where $\mathbf{R}$ is the residual force vector. The internal nodal force vector is evaluated by integrating the stresses over the element domain as
$\mathbf{F}_{\mathrm{int}}^{e}=\int_{\Omega^{e}}\overset{\sim}{\mathbf{B}}^{\mathrm{T}}\boldsymbol{\sigma}\,\mathrm{d}\Omega.$ (26)
During each iteration, the stress increment is obtained from
$\Delta\boldsymbol{\sigma}=\mathbf{D}_{ep}\Delta\boldsymbol{\varepsilon},$ (27)
and the residual force is updated accordingly. The nonlinear system is solved using a Newton–Raphson scheme:
$\mathbf{K}_{t}^{(i)}\Delta\mathbf{u}^{(i)}=\mathbf{R}^{(i)},$ (28)
where $\mathbf{K}_{t}^{(i)}$ is the current tangent stiffness matrix. Iterations proceed until the normalized residual force satisfies
$\frac{\|\mathbf{R}^{(i)}\|}{\|\mathbf{F}_{\mathrm{ext}}\|}\leq\varepsilon_{\mathrm{tol}}.$ (29)
This residual force evaluation ensures equilibrium at each load step and enables accurate simulation of nonlinear elasto-plastic deformation.